-70=D/10 =D
P/12.4=3/2 P=
23=X-15 =X
2X=62 =X

Answer by checkley75(3666)

-70=D/10 CROSS MULTIPLY.
D=10*70
D=700
---------------------------
P/12.4=3/2 CROSS MULTIPLY
2P=3*12.4
2P=37.2 CROSS MULTIPLY.
P=37.2/2
P=18.6 ANSWER.
PROOF
18.6/12.4=3/2
37.2=37.2